# Graph Optimality-Aware Stochastic LiDAR Bundle Adjustment with Progressive Spatial Smoothing

### v1

@@ 2.3 Optimality of LiDAR Bundle Adjustment
where $\mathbf{J}^{\epsilon}_{k}$ is the Jacobian matrix of $\bm{\epsilon}_{k}$ with respect to $\mathbf{X}$. The information matrix $\bm{\varLambda}$ reveals the reliability of the poses to be estimated in bundle adjustment and also the optimality of the graph $\mathbf{G}$. The optimality of bundle adjustment for $\mathbf{G}$ is defined using the minimal eigenvalue of the information matrix: $\lambda_{min}(\bm{\varLambda})$ (Khanna et al., 2017; Pukelsheim, 2006).
2.4 Schur Complement for State Marginalization
@@
It is desirable that if $\mathbf{H}_{bb}$ is sparse for the inverse operation. The sparsity of $\mathbf{H}_{bb}$ can be achieved by using edge selection, which will be detailed in Section 5. Given an optimal update $\Delta\mathbf{X}_{a}$, the update for the remaining $\Delta\mathbf{X}_{b}$ is calculated by (12).
@@
3 System Overview
To achieve a robust, efficient, and scalable LBA, the proposed method shown in Fig.2 includes two key modules, progressive spatial smoothing (PSS) detailed in Section 4 and graph optimality-aware stochastic optimization (GOSO) detailed in Section 5.
Taking the inaccurate initial LiDAR frames as input, GOSO first determines the relation graph of the LiDAR frames so that it achieves efficient optimization (Section 5.1) according to the optimality of the graph. Second, PSS extracts the factors from the remaining edges in the sparse relation graph. It utilizes a spatial smoothing kernel to extract the LiDAR correspondences and a polynomial surface function to construct the factors for robust pose correction. Then, GOSO separates the raw graph into subsets according to modularity (Blondel et al., 2008) considering stochastic condition to avoid local minimal in the following optimization while maintaining good relationships in each subset (Section 5.2). Finally, GOSO optimizes the LiDAR frames with graph marginalization in a successive manner to achieve a scalable optimization (Section 5.3) for large-scale LBA. The above steps are repeated until the optimization converges or reaches the preset max iteration times (5 used in the experiment). After each iteration, the spatial kernel size $\gamma$ is decreased by $\gamma/T_{D}$. The detailed descriptions of the proposed method are as follows.
4 Progressive Spatial Smoothing (PSS)
Extracting reliable LiDAR correspondences from LiDAR frames and constructing appropriate types of factors are crucial for robust pose adjustment. Noting that structural prior information can still be obtained even with imperfect initial poses, the proposed Progressive Spatial Smoothing (PSS) first extracts LiDAR correspondences and structural prior information through spatial smoothing. Subsequently, polynomial surface factors are constructed for the bundle adjustment process. It is important to note that only the correspondences between the remaining edges in the sparse relation graph (Section 5.1) are considered to ensure efficiency.
4.1 Spatial Smoothing for LiDAR Correspondences
4.1.1 Smoothing Kernel Sampling
@@
where $\mathbf{D}(\hat{\mathbf{n}}_{i})$ is the differential function for $\hat{\mathbf{n}}_{i}$ in the 3D space. The $L_{0}$ normalization counting the non-zero item is used here to eliminate the influence of outliers and preserve the original shape in the sharp regions (Xu et al., 2011; Sun et al., 2015). $\mu$ is the weight that balances the data term and smooth term. Equation (14) can be minimized using an auxiliary function which is detailed in A.
@@
4.1.2 Weighted Surface Fitting
@@
Now the core problem is to robustly determine the optimal parameter $\bm{\alpha}^{i}$ using the projected noisy points $\{\hat{\mathbf{p}}^{S}_{j},j\in\rm{\Phi}_{i}\}$. We find the best parameters $\bm{\alpha}^{i}$ using least-square estimation considering the Gaussian radial weight function $\mathrm{w}(d)$ (Alexa et al., 2003):
@@
4.1.3 Points Smoothing and Factor Association
@@
4.2 Polynomial Surface Factor
@@
$\sigma_{i,j}$ is correlated with $i^{th}$ and $j^{th}$ poses, namely $[\hat{R}_{i},\hat{t}_{i}]$ and $[\hat{R}_{j},\hat{t}_{j}]$ as shown in Fig. 3. The Jacobian of $\sigma_{i,j}$ with respect to the $i^{th}$ pose can be derived using the chain rule in C.
5 Graph Optimalily-Aware Stochastic Optimization (GOSO)
5.1 Optimality-aware Graph Sparsification
To build the relation graph $\mathbf{G}=(\mathbf{X},\mathbf{E})$ for the LBA problem, the LiDAR frames are first projected to the world frame according to the initial poses $\mathbf{\breve{X}}$. The projected point clouds are then voxelized using the voxel size of $\gamma$. In each voxel, it records the points from different frames. Any two frames that share more than $30\%$ overlaps of voxels will be used to establish an edge in $\mathbf{G}$. Meanwhile, the reliability of the edge between two frames will be calculated according to pair-wise point cloud registration (B).
The relation graph $\mathbf{G}$ usually contains dense edges due to the complex overlaps between the frames, making the solving procedure of bundle adjustment have low-efficiency. Thus, we sparsify $\mathbf{G}$ considering the optimal, namely, the minimal eigenvalue of the information matrix $\lambda_{min}(\bm{\varLambda})$ mentioned in Section 2.3. To make the raw dense relation graph $\mathbf{G}$ more sparse by only maintaining a subset of the raw edges $\mathbf{E}$ while keeping good optimality, it is formulated as an optimal design problem. To achieve this optimal design, we try to optimize the following objective function (20).
$\begin{split}\mathrm{argmax}\leavevmode\nobreak\ \mathrm{F}_{set}(\mathbf{E}^{o})\leavevmode\nobreak\ &s.t.\leavevmode\nobreak\ \mathbf{E}^{o}\subset\mathbf{E},|\mathbf{E}^{o}|\leqslant\mathbf{N}^{o},\\ \mathrm{F}_{set}(\circ)&=\lambda_{min}(\bm{\varLambda}(\circ)),\end{split}$ (20)
where $\mathbf{N}^{o}$ is the desired number of remaining edges $\mathbf{E}^{o}$ in the optimal designed graph $\mathbf{G}^{o}=(\mathbf{X},\mathbf{E}^{o})$. Equation (20) is an NP-hard problem, while it can still be approximately solved efficiently using stochastic-greedy heuristic searching (Khanna et al., 2017). The stochastic-greedy algorithm starts with an empty set. Then, at each step, it selects one element from a random set with a size of $\mathrm{log}(1/\epsilon)|\mathbf{E}|$ from the remaining edge set, which gains most of the objective function $\mathrm{F}_{set}(\mathbf{E}^{o})$. $\epsilon$ is the decay factor. The iterative process stops until it gets $\mathbf{N}^{o}$ edges. In the real experiment, $\mathbf{N}^{o}$ can be set according to a certain percentage (e.g., $20\%$) of the size of raw edges. The details of the stochastic-greedy heuristic searching can be found in algorithm 1.
@@ 9       $\bm{\varLambda}(\mathbf{E}^{o})\leftarrow\bm{\varLambda}(\mathbf{E}^{o})+\bm{\varLambda}_{v*}$;
5.2 Stochastic Graph Clustering
@@
The proposed stochastic graph clustering operates on the relation graph $\mathbf{G}$, where the weight of each edge $\mathbf{E}_{k}$ is measured using the minimal eigenvalue of the covariance $\lambda^{min}_{k}$, which is introduced in Section 2.3. At the beginning, each node formulates an individual cluster. Next, the nodes that share an edge will be merged as a new cluster. In order to make the nodes in each cluster share a tight relationship, an optimal clustering is to be found to maximize the modularity $\mathbf{Q}$ (22(Blondel et al., 2008).
@@
Finding the optimal clustering that maximizes the modularity is an NP-hard problem. Louvain’s algorithm (Blondel et al., 2008), greedily merging two clusters that give the maximum increase of the modularity is a standard and efficient strategy to solve the problem (Blondel et al., 2008). However, Louvain’s algorithm lacks randomization due to the greedy strategy. Inspired by Zhou et al. (2020), we randomly merge clusters according to a probability distribution based on modularity increments. In other words, the larger the modularity increment, the more likely the two clusters are to merge. The maximum size of the items in each cluster is set to $T_{\mathrm{cluster}}$ (e.g., 300 is used in the experiment). The random merging is terminated when the modularity is not increasing. Figure 4 illustrates the different clustering results of a relation graph. The proposed stochastic graph clustering achieves different but meaningful results in each iteration to achieve high accuracy while avoiding local minimum.
5.3 Graph Marginalization
@@
Dividing the raw large-scale bundle adjustment problem into subproblems is a common strategy for the image or LBA. For example, the hierarchical LBA (Liu et al., 2023) split the raw problem into subproblems according to the length of the successive frames, which does not consider the spatial relationship between the frames. Our problem-dividing strategy is inspired by the large-scale image bundle adjustment (Jiang et al., 2024) and the reduced camera system (Demmel et al., 2021), which consider the spatial relationships between the images using the initial values of the cameras leading to better accuracy of the final results.
5.3.1 Extra-Cluster Graph Optimization
@@
5.3.2 Intra-Cluster Adjustment with Marginalization
@@ 6.1 Implementation
6.2 Evaluation on Ground-based Benchmark Dataset
We first evaluate the proposed PSS-GOSO on the ground-based benchmark dataset, named FusionPortable (Wei et al., 2024). The FusionPortable contained diverse platforms and diverse environments, covering the main scenes that need accurate 3D mapping. Using the initial poses obtained by Fast-lio (Xu et al., 2022), the dense graph and point clouds generated by PSS-GOSO in the diverse environments are demonstrated in Fig. 6. We mainly focuse on the handheld, legged, and UGV platforms which often operate in complex environments without GNSS signals. We compare the proposed method with the existing pose-graph-based method (Koide et al., 2019). The APE distribution of the proposed method in the handheld, legged, and UGV sequences are plotted in Fig. 7, Fig. 8, and Fig. 9, respectively. The APE of different methods are listed in Table. 1, which demonstrate that the proposed PSS-GOSO improved the initial pose accuracies. However, the existing pose-graph-based method (Koide et al., 2019) may suffer from inaccurate pair-wise registration thus achieving less accuracy.
Table 1: Absolute Position Error (APE) of different methods in FusionPortable (Wei et al., 2024) dataset [m]. The best results are in BOLD.
Sequence Initial  (Xu et al., 2022) PSS-GOSO Pose-graph (Koide et al., 2019) HBA Liu et al. (2023)
@@ 6.2 Evaluation on Ground-based Benchmark Dataset
Figure 6: The point clouds and dense graphs in FusionPortable (Wei et al., 2024) dataset. Edges of the graph are marked as yellow lines. The nodes of the graph are marked as green axes.
Figure 7: Absolute Position Error (APE) distribution of PSS-GOSO in the handheld sequences of FusionPortable (Wei et al., 2024) dataset.
Figure 8: Absolute Position Error (APE) distribution of PSS-GOSO in the legged sequences of FusionPortable (Wei et al., 2024) dataset.
Figure 9: Absolute Position Error (APE) distribution of PSS-GOSO in the UGV sequences of FusionPortable (Wei et al., 2024) dataset.
6.3 Evaluation on Air-based Benchmark Dataset
Aerial LiDAR point cloud quality is crucial for photogrammetry, especially in GNSS-denied environments. We assess the performance of our proposed PSS-GOSO using the MARS-LVIG benchmark dataset (Li et al., 2024a), which includes diverse environments such as islands, valleys, airports, and towns. Initial pose estimates were obtained using Fast-LIO (Xu et al., 2022). The dense graph and point clouds generated by PSS-GOSO in the diverse environments are demonstrated in Fig. 10. The APE distribution of the proposed method in the UAV sequences is plotted in Fig. 11. The APE of different methods is listed in Table. 2. The results demonstrate that PSS-GOSO significantly improves accuracy compared to other methods. For instance, in the island sequence, PSS-GOSO reduces the APE to 1.14 meters, improving on the initial 2.37 meters and surpassing both the pose-graph-based method (Koide et al., 2019) and the HBA method (Liu et al., 2023). In the valley sequence, PSS-GOSO achieves an APE of 4.75 meters, a substantial reduction from the initial 11.99 meters and better than the pose-graph method’s 7.37 meters and HBA’s 6.40 meters. For the airport sequence, PSS-GOSO’s APE of 1.60 meters outperforms the initial 3.06 meters and the results from other methods. In the town sequence, PSS-GOSO achieves a notable APE improvement compared to the initial 2.37 meters, with better results than the pose-graph method and HBA. These results highlight PSS-GOSO’s effectiveness in enhancing the precision of aerial LiDAR point clouds across various challenging environments.
Table 2: Absolute Position Error (APE) of different methods in MARS-LVIG (Li et al., 2024a) dataset [m]. The best results are in BOLD.
Sequence Initial (Xu et al., 2022) PSS-GOSO Pose-graph  (Koide et al., 2019) HBA (Liu et al., 2023)

### v2

@@ II-C Optimality of LiDAR Bundle Adjustment
where $\mathbf{J}^{\epsilon}_{k}$ is the Jacobian matrix of $\bm{\epsilon}_{k}$ with respect to $\mathbf{X}$. The information matrix $\bm{\varLambda}$ reveals the reliability of the poses to be estimated in bundle adjustment and also the optimality of the graph $\mathbf{G}$. The optimality of bundle adjustment for $\mathbf{G}$ is defined using the minimal eigenvalue of the information matrix: $\lambda_{min}(\bm{\varLambda})$ [26].
II-D Schur Complement for State Marginalization
@@
It is desirable that if $\mathbf{H}_{bb}$ is sparse for the inverse operation. The sparsity of $\mathbf{H}_{bb}$ can be achieved by using edge selection, which will be detailed in Section V. Given an optimal update $\Delta\mathbf{X}_{a}$, the update for the remaining $\Delta\mathbf{X}_{b}$ is calculated by (12).
@@
III System Overview
To achieve a robust, efficient, and scalable LBA, the proposed method shown in Fig.2 includes two key modules, progressive spatial smoothing (PSS) detailed in Section IV and graph optimality-aware stochastic optimization (GOSO) detailed in Section V.
Taking the inaccurate initial LiDAR frames as input, GOSO first determines the relation graph of the LiDAR frames so that it achieves efficient optimization (Section V-A) according to the optimality of the graph. Second, PSS extracts the factors from the remaining edges in the sparse relation graph. It utilizes a spatial smoothing kernel to extract the LiDAR correspondences and a polynomial surface function to construct the factors for robust pose correction. Then, GOSO separates the raw graph into subsets according to modularity [27] considering stochastic condition to avoid local minimal in the following optimization while maintaining good relationships in each subset (Section V-B). Finally, GOSO optimizes the LiDAR frames with graph marginalization in a successive manner to achieve a scalable optimization (Section V-C) for large-scale LBA. The above steps are repeated until the optimization converges or reaches the preset max iteration times (5 used in the experiment). After each iteration, the spatial kernel size $\gamma$ is decreased by $\gamma/T_{D}$. The detailed descriptions of the proposed method are as follows.
IV Progressive Spatial Smoothing (PSS)
Extracting reliable LiDAR correspondences from LiDAR frames and constructing appropriate types of factors are crucial for robust pose adjustment. Noting that structural prior information can still be obtained even with imperfect initial poses, the proposed Progressive Spatial Smoothing (PSS) first extracts LiDAR correspondences and structural prior information through spatial smoothing. Subsequently, polynomial surface factors are constructed for the bundle adjustment process. It is important to note that only the correspondences between the remaining edges in the sparse relation graph (Section V-A) are considered to ensure efficiency.
IV-A Spatial Smoothing for LiDAR Correspondences
IV-A1 Smoothing Kernel Sampling
@@
where $\mathbf{D}(\hat{\mathbf{n}}_{i})$ is the differential function for $\hat{\mathbf{n}}_{i}$ in the 3D space. The $L_{0}$ normalization counting the non-zero item is used here to eliminate the influence of outliers and preserve the original shape in the sharp regions [28, 29]. $\mu$ is the weight that balances the data term and smooth term. Equation (14) can be minimized using an auxiliary function which is detailed in Appendix A: Auxiliary optimization.
@@
IV-A2 Weighted Surface Fitting
@@
Now the core problem is to robustly determine the optimal parameter $\bm{\alpha}^{i}$ using the projected noisy points $\{\hat{\mathbf{p}}^{S}_{j},j\in\rm{\Phi}_{i}\}$. We find the best parameters $\bm{\alpha}^{i}$ using least-square estimation considering the Gaussian radial weight function $\mathrm{w}(d)$ [30]:
@@
IV-A3 Points Smoothing and Factor Association
@@
IV-B Polynomial Surface Factor
@@
$\sigma_{i,j}$ is correlated with $i^{th}$ and $j^{th}$ poses, namely $[\hat{R}_{i},\hat{t}_{i}]$ and $[\hat{R}_{j},\hat{t}_{j}]$ as shown in Fig. 3. The Jacobian of $\sigma_{i,j}$ with respect to the $i^{th}$ pose can be derived using the chain rule in Appendix C: Jacobian of Progressive Spatial Smoothing (PSS) Factor.
V Graph Optimalily-Aware Stochastic Optimization (GOSO)
V-A Optimality-aware Graph Sparsification
To build the relation graph $\mathbf{G}=(\mathbf{X},\mathbf{E})$ for the LBA problem, the LiDAR frames are first projected to the world frame according to the initial poses $\mathbf{\breve{X}}$. The projected point clouds are then voxelized using the voxel size of $\gamma$. In each voxel, it records the points from different frames. Any two frames that share more than $30\%$ overlaps of voxels will be used to establish an edge in $\mathbf{G}$. Meanwhile, the reliability of the edge between two frames will be calculated according to pair-wise point cloud registration (Appendix B: Relative Pose Residual and Covariance from Pair-wise Point Cloud Registration).
The relation graph $\mathbf{G}$ usually contains dense edges due to the complex overlaps between the frames, making the solving procedure of bundle adjustment have low-efficiency. Thus, we sparsify $\mathbf{G}$ considering the optimal, namely, the minimal eigenvalue of the information matrix $\lambda_{min}(\bm{\varLambda})$ mentioned in Section II-C. To make the raw dense relation graph $\mathbf{G}$ more sparse by only maintaining a subset of the raw edges $\mathbf{E}$ while keeping good optimality, it is formulated as an optimal design problem. To achieve this optimal design, we try to optimize the following objective function (20).
$\begin{split}\mathrm{argmax}~{}\mathrm{F}_{set}(\mathbf{E}^{o})~{}&s.t.~{}\mathbf{E}^{o}\subset\mathbf{E},|\mathbf{E}^{o}|\leqslant\mathbf{N}^{o},\\ \mathrm{F}_{set}(\circ)&=\lambda_{min}(\bm{\varLambda}(\circ)),\end{split}$ (20)
where $\mathbf{N}^{o}$ is the desired number of remaining edges $\mathbf{E}^{o}$ in the optimal designed graph $\mathbf{G}^{o}=(\mathbf{X},\mathbf{E}^{o})$. Equation (20) is an NP-hard problem, while it can still be approximately solved efficiently using stochastic-greedy heuristic searching [26]. The stochastic-greedy algorithm starts with an empty set. Then, at each step, it selects one element from a random set with a size of $\mathrm{log}(1/\epsilon)|\mathbf{E}|$ from the remaining edge set, which gains most of the objective function $\mathrm{F}_{set}(\mathbf{E}^{o})$. $\epsilon$ is the decay factor. The iterative process stops until it gets $\mathbf{N}^{o}$ edges. In the real experiment, $\mathbf{N}^{o}$ can be set according to a certain percentage (e.g., $20\%$) of the size of raw edges. The details of the stochastic-greedy heuristic searching can be found in algorithm 1.
@@ 9       $\bm{\varLambda}(\mathbf{E}^{o})\leftarrow\bm{\varLambda}(\mathbf{E}^{o})+\bm{\varLambda}_{v*}$;
V-B Stochastic Graph Clustering
@@
The proposed stochastic graph clustering operates on the relation graph $\mathbf{G}$, where the weight of each edge $\mathbf{E}_{k}$ is measured using the minimal eigenvalue of the covariance $\lambda^{min}_{k}$, which is introduced in Section II-C. At the beginning, each node formulates an individual cluster. Next, the nodes that share an edge will be merged as a new cluster. In order to make the nodes in each cluster share a tight relationship, an optimal clustering is to be found to maximize the modularity $\mathbf{Q}$ (22[27].
@@
Finding the optimal clustering that maximizes the modularity is an NP-hard problem. Louvain’s algorithm [27], greedily merging two clusters that give the maximum increase of the modularity is a standard and efficient strategy to solve the problem [27]. However, Louvain’s algorithm lacks randomization due to the greedy strategy. Inspired by [33], we randomly merge clusters according to a probability distribution based on modularity increments. In other words, the larger the modularity increment, the more likely the two clusters are to merge. The maximum size of the items in each cluster is set to $T_{\mathrm{cluster}}$ (e.g., 300 is used in the experiment). The random merging is terminated when the modularity is not increasing. Figure 4 illustrates the different clustering results of a relation graph. The proposed stochastic graph clustering achieves different but meaningful results in each iteration to achieve high accuracy while avoiding local minimum.
V-C Graph Marginalization
@@
Dividing the raw large-scale bundle adjustment problem into subproblems is a common strategy for the image or LBA. For example, the hierarchical LBA [2] split the raw problem into subproblems according to the length of the successive frames, which does not consider the spatial relationship between the frames. Our problem-dividing strategy is inspired by the large-scale image bundle adjustment [34] and the reduced camera system [35], which consider the spatial relationships between the images using the initial values of the cameras leading to better accuracy of the final results.
V-C1 Extra-Cluster Graph Optimization
@@
V-C2 Intra-Cluster Adjustment with Marginalization
@@ VI-A Implementation
VI-B Evaluation on Ground-based Benchmark Dataset
We first evaluate the proposed PSS-GOSO on the ground-based benchmark dataset, named FusionPortable [36]. The FusionPortable contained diverse platforms and diverse environments, covering the main scenes that need accurate 3D mapping. Using the initial poses obtained by Fast-lio [37] and manual loop closure, the dense graph and point clouds generated by PSS-GOSO in the diverse environments are demonstrated in Fig. 6. We mainly focus on the handheld, legged, and UGV platforms which often operate in complex environments without GNSS. We compare the proposed method with the existing pose-graph-based method [38]. The APE distribution of the proposed method in the handheld, legged, and UGV sequences are plotted in Fig. 7, Fig. 8, and Fig. 9, respectively. The APE of different methods are listed in Table. I, which demonstrate that the proposed PSS-GOSO improved the initial pose accuracies. However, the existing pose-graph-based method  [38] may suffer from inaccurate pair-wise registration thus achieving less accuracy.
TABLE I: Absolute Position Error (APE) of different methods in FusionPortable [36] dataset [m]. The best results are in BOLD.
Sequence Initial (SLAM + Loop)  [37] PSS-GOSO Pose-graph [38] HBA [2]
@@ VI-B Evaluation on Ground-based Benchmark Dataset
Figure 6: The point clouds and dense graphs in FusionPortable [36] dataset. Edges of the graph are marked as yellow lines. The nodes of the graph are marked as green axes.
Figure 7: Absolute Position Error (APE) distribution of PSS-GOSO in the handheld sequences of FusionPortable [36] dataset.
Figure 8: Absolute Position Error (APE) distribution of PSS-GOSO in the legged sequences of FusionPortable [36] dataset.
Figure 9: Absolute Position Error (APE) distribution of PSS-GOSO in the UGV sequences of FusionPortable [36] dataset.
VI-C Evaluation on Air-based Benchmark Dataset
Aerial LiDAR point cloud quality is crucial for photogrammetry. However, GNSS is very vulnerable to accidental or malicious radio frequency interruptions (jamming) or fake signals (spoofing) [39, 40]. We assess the performance of our proposed PSS-GOSO without GNSS singals using the MARS-LVIG benchmark dataset [41], which includes diverse environments such as islands, valleys, airports, and towns. Initial pose estimates were obtained using Fast-LIO [37] and manual loop closure. The dense graph and point clouds generated by PSS-GOSO in the diverse environments are demonstrated in Fig. 10. The APE distribution of the proposed method in the UAV sequences is plotted in Fig. 11. The APE of different methods is listed in Table. II. The results demonstrate that PSS-GOSO significantly improves accuracy compared to other methods. For instance, in the island sequence, PSS-GOSO reduces the APE to 1.14 meters, improving on the initial 2.37 meters and surpassing both the pose-graph-based method [38] and the HBA method [2]. In the valley sequence, PSS-GOSO achieves an APE of 4.75 meters, a substantial reduction from the initial 11.99 meters and better than the pose-graph method’s 7.37 meters and HBA’s 6.40 meters. For the airport sequence, PSS-GOSO’s APE of 1.60 meters outperforms the initial 3.06 meters and the results from other methods. In the town sequence, PSS-GOSO achieves a notable APE improvement compared to the initial 2.37 meters, with better results than the pose-graph method and HBA. These results highlight PSS-GOSO’s effectiveness in enhancing the precision of aerial LiDAR point clouds across various challenging environments.
TABLE II: Absolute Position Error (APE) of different methods in MARS-LVIG [41] dataset [m]. The best results are in BOLD.
Sequence Initial (SLAM + Loop)[37] PSS-GOSO Pose-graph  [38] HBA [2]